f(x)=x^2-4x+3.
what is the domain and range?
determine if it is odd,even or neither.
: f(x)=x^2-4x+3.
what is the domain and range?
determine if it is odd,even or neither.

Answer by edjones(2391)

ax^2+bx+c
x^2-4x+3
(x-3)(x-1)=0
x=3, x=1 when y=0
The graph of the parabola has a minimum because a is positive.
(3+1)/2=2 so y will be at a minimum when x=2
f(x)=y=2^2-(4*2)+3
=4-8+3
=-1
The domain of x is all real numbers.
The range of y is -1 to infinity.
It is neither odd nor even with respect to the axis of the graph.
.
Ed
.
graph(500,500,-10,10,-10,10,x^2-4x+3)